OpenStudy (anonymous):

the value |z|, if z satisfies the equation a(z+a^2)=2b-z, where a =-3.679+(-2.699)i and b = 0.474 + (-0.743)i, is: .............?

OpenStudy (anonymous):

where i = \[\sqrt{-1}\]

OpenStudy (dumbcow):

\[\large z = \frac{2b-a^{3}}{a+1} = \frac{-29.657 +88.446i}{-2.679-2.699i}\] \[\large z = \frac{-159.265 -316.991i}{14.46}\] \[\large z = -11.014 -21.92i\] \[\large |z| = \sqrt{601.79} = 24.53\]

OpenStudy (anonymous):

thanks! perfect, it's been a while since i've done imaginary numbers

OpenStudy (dumbcow):

yeah these numbers are kinda tedious to do by hand...i cheated with wolfram

OpenStudy (dumbcow):

Basically remember that: if z = a+bi |z| = sqrt[(a+bi)(a-bi)]

OpenStudy (anonymous):

okay noted, thanks again. love wolfram